OpenStudy (anonymous):

Due in 4 minutes! Help! I just need answer please A movie star, unwilling to give his age, posed the following riddle to a gossip columnist: "6 years ago, I was 13 times as old as my daughter. Now I am 7 times as old as she is." How old is the movie star?

OpenStudy (anonymous):

Age of the star now = a Age of the daughter now = d a - 6 = 13 (d - 6) a = 7 * d a - 6 = 13d - 78 a = 13d - 72 13d - 72 = 7d 6d = 72 d = 12

OpenStudy (anonymous):

so her age is 12?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

or 7 times 12?

OpenStudy (anonymous):

7 times twelve is the age of her father, the movie star.